OpenStudy (anonymous):

find the particular antiderivative that satisfies the following conditions: p(prime)(x)=70/(x^3) and p(2)=4

OpenStudy (amistre64):

\[70\int x^{-3}dx \]

OpenStudy (anonymous):

but then p(2) is not giving me 4

OpenStudy (amistre64):

\[F(x)=-\frac{35}{x^2}\]

OpenStudy (amistre64):

+C

OpenStudy (anonymous):

the antiderivative is \[P(x)=-35x^{-2}+c\] now put \[P(2)=-\frac{35}{2^2}+c=4\] and solve for c

OpenStudy (amistre64):

yeps; this thing is slooowwww today.

OpenStudy (anonymous):

slow slow slow. i don 't even see the latex i wrote

OpenStudy (anonymous):

lol oh I see thanks!

OpenStudy (anonymous):

did you get that? you solve for the constant by the information given. i can't seem to write it

OpenStudy (anonymous):

\[P(x)=-35x^{-2}+c\] now put \[P(2)=-\frac{35}{2^2}+c=4\] and solve for c

OpenStudy (anonymous):

ok forget it. i'm out